# Prediction-Powered Adaptive Shrinkage Estimation

### v2

@@ 1 Introduction
A major obstacle in answering modern scientific questions is the scarcity of gold-standard data (Miao et al., 2024b). While advancements in data collection, such as large-scale astronomical surveys (York et al., 2000) and web crawling (Penedo et al., 2024), have led to an abundance of covariates (or features), scientific conclusions often rely on outcomes (or labels), which are often expensive and labor-intensive to obtain. The rapid development of machine learning (ML) algorithms has offered a path forward, with ML predictions increasingly used to supplement gold-standard outcomes and increase the statistical efficiency of subsequent analyses (Liang et al., 2007; Wang et al., 2020).
Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023) addresses the scarcity issue by providing a framework for valid statistical analysis using predictions from black-box ML models. By combining ML-predicted and gold-standard outcomes, PPI and its variants (Angelopoulos et al., 2024; Zrnic & Candès, 2024; Zrnic, 2025) use the abundance of predictions to reduce variance while relying on the accuracy of labeled [^1] Throughout the paper, we use the terms “labeled” and “gold-standard” interchangeably. data to control bias.
In this work, we adapt PPI to the estimation of multiple outcome means in compound estimation settings. Many applications of PPI naturally involve parallel statistical problems that can be solved simultaneously. For instance, several PPI methods (Angelopoulos et al., 2024; Fisch et al., 2024) have shown improvements in estimating the fraction of spiral galaxies using predictions on images from the Galaxy Zoo 2 dataset (Willett et al., 2013). While these methods focus on estimating a single overall fraction, a richer analysis emerges from partitioning galaxies based on metadata (such as celestial coordinates or pre-defined bins) and estimating the fraction of galaxies within each partition. This compound estimation approach enables more granular scientific inquiries that account for heterogeneity across galaxy clusters and spatial locations (Nair & Abraham, 2010).
We demonstrate, both theoretically and empirically, the benefits of solving multiple mean estimation problems simultaneously. Our approach builds on the empirical Bayes (EB) principle of sharing information across problems (Robbins, 1956; Efron, 2010) as exemplified by James-Stein shrinkage (James & Stein, 1961; Xie et al., 2012). The connection between modern and classical statistical ideas allows us to perform within problem PPI estimation in the first place, followed by a shrinkage step reusing the ML predictions in an adaptive way, which becomes possible through borrowing information across problems. Our contributions are:
@@ 2.1 Prediction-Powered Inference (PPI)
The PPI framework considers a setting where we have access to a small number of labeled data points $(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a large number of unlabeled covariates $(\tilde{X}_{i})_{i=1}^{N}\in(\mathcal{X})^{N}$, where $\mathcal{X}$ and $\mathcal{Y}$ represent the covariate and outcome space, respectively. The data points are drawn iid from a joint distribution $\mathbb{P}_{XY}$. [^2] To be concrete: $(X_{i},Y_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$ and $(\tilde{X}_{i},\tilde{Y}_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$ independently, but $\tilde{Y}_{i}$ is unobserved. We are also given a black-box predictive model $f:\mathcal{X}\to\mathcal{Y}$ that is independent of the datasets (e.g., pre-trained on similar but unseen data). For mean estimation with $\mathcal{Y}\subset\mathbb{R}$, the goal is to leverage the predicted outcomes $f(X_{i})$ to improve the estimation of $\theta:=\mathbb{E}[Y_{i}]$. Some simple estimators take the form of the following aggregated (summary) statistics
$\bar{Y}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\>\,$ ${\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}\tilde{Y}:=\frac{1}{N}\sum_{i=1}^{N}\tilde{Y}_{i},}$ (1)
$\bar{Z}^{f}:=\frac{1}{n}\sum_{i=1}^{n}f(X_{i}),\>\,$ $\tilde{Z}^{f}:=\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i}).$
@@
Both definitions represent $\hat{\theta}^{\text{PPI}}$ in the form of a baseline estimator plus a correction term. In the first representation, the baseline estimator is the unbiased classical estimator $\bar{Y}$, while the correction term has expectation 00 and attempts to reduce the variance of $\bar{Y}$. In the second representation, the baseline estimator is the prediction mean on unlabeled data $\tilde{Z}^{f}$ (which in general may be biased for $\theta$), while the correction term removes the bias of $\tilde{Z}^{f}$ by estimating the bias of the ML model $f$ on the labeled dataset. Writing $\hat{\theta}^{\text{PPI}}=\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i})+\tfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))$, we find that $\mathbb{E}[\hat{\theta}^{\text{PPI}}]=\mathbb{E}[Y_{i}]=\theta$ and
@@
To overcome the above limitation, Angelopoulos et al. (2024) introduce a power-tuning parameter $\lambda$ and define
@@
Consider the ideal scenario for PPI with $N=\infty$ (that is, the unlabeled dataset is much larger than the labeled dataset) so that $\tilde{Z}^{f}\equiv\mathbb{E}[f(\tilde{X}_{i})]$. Even then, the MSE of $\hat{\theta}^{\text{PPI}}$ in (3) is always lower bounded [^5] The same lower bound also applies to power-tuned PPI $\hat{\theta}^{\text{PT}}$. by $\tfrac{1}{n}\mathbb{E}[\mathrm{Var}[Y_{i}\mid X_{i}]]$ and the lower bound is attained by the perfect ML predictor $f(\cdot)\equiv\mathbb{E}\!\left[Y_{i}\mid X_{i}=\cdot\right]$. In words, if $Y_{i}$ is not perfectly predictable from $X_{i}$, then PPI applied to a labeled dataset of fixed size $n$ must have non-negligible MSE. By contrast, for $N=\infty$, the prediction mean of unlabeled data $\tilde{Z}^{f}$ has zero variance and MSE equal to the squared bias $(\mathbb{E}[f(X_{i})]-\theta_{i})^{2}$. Thus if the predictor satisfies a calibration-type property that $\mathbb{E}[f(X_{i})]\approx\mathbb{E}[Y_{i}]$ (which is implied by, but much weaker than the requirement $f(X_{i})\approx Y_{i}$), then the MSE of $\tilde{Z}^{f}$ could be nearly 00. By contrast, PPI (and PPI++) can only partially capitalize on such a predictor $f(\cdot)$.
@@ 2.2 The Compound Mean Estimation Setting
In this section, we introduce the problem setting that PAS is designed to address—estimating the mean of $m>1$ parallel problems with a single black-box predictive model $f$ [^6] Our proposal also accommodates using separate predictors $\{f_{j}\}_{j=1}^{m}$ for each problem. To streamline exposition, we focus on the practical scenario where a single (large) model (e.g., a large language or vision model) can handle multiple tasks simultaneously (Radford et al., 2019; He et al., 2022). For the $j$-th problem, where $j\in[m]:=\{1,\ldots,m\}$, we observe a labeled dataset $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ samples and an unlabeled dataset $(\tilde{X}_{ij})_{i=1}^{N_{j}}$ with $N_{j}\in\mathbb{N}$ samples. We start with modeling heterogeneity across problems.
@@
$\eta_{j}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{\eta},\;j\in[m],\,\,\text{and}\,\,\boldsymbol{\eta}:=(\eta_{1},...,\eta_{m})^{\intercal},$ (5)
@@
We do not place any restriction over the unknown prior $\mathbb{P}_{\eta}$. Assumption 2.1 posits exchangeability across problems, which enables information sharing, without restricting heterogeneity (Ignatiadis et al., 2023). In our setting, we are specifically interested in the means
@@
$\begin{bmatrix}f(X_{ij})\\ Y_{ij}\end{bmatrix}\,\big{|}\,\eta_{j}\,\,{\buildrel\text{iid}\over{\sim}\,}\,\mathbb{F}_{j}\left(\begin{bmatrix}\mu_{j}\\ \theta_{j}\end{bmatrix},\begin{bmatrix}\tau_{j}^{2}&\rho_{j}\tau_{j}\sigma_{j}\\ \rho_{j}\tau_{j}\sigma_{j}&\sigma_{j}^{2}\end{bmatrix}\right),$ (7)
@@
Similar to Eq. (1), we define the aggregated statistics $\bar{Y}_{j},\bar{Z}_{j}^{f},\tilde{Z}^{f}_{j}$ for each $j\in[m]$. To facilitate exposition, following prior work, [^7] For EB, examples include Xie et al. (2012); Soloff et al. (2024); for PPI, see recent works like Fisch et al. (2024). we treat the second moments $\tau_{j}^{2}$, $\sigma_{j}^{2}$, $\gamma_{j}$ as known until the end of Section 5. In Section 6, we extend our methods to allow for unknown $\tau_{j}^{2}$, $\sigma_{j}^{2}$, and $\gamma_{j}$.
@@
$X_{ij}\,{\buildrel\text{iid}\over{\sim}\,}\mathcal{N}(\eta_{j},\psi^{2}),\;\;\;Y_{ij}|X_{ij}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(2\eta_{j}X_{ij}-\eta_{j}^{2},c).$ (8)
@@
To conclude this section, we define the compound risk (Robbins, 1951; Jiang & Zhang, 2009) for any estimator $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{m})^{\intercal}$ as the expected squared error loss averaged over problems,
@@ 3 Statistical Guiding Principles & Prior Work
An important statistical idea is to improve $\hat{\theta}^{\text{cl}}$ via obtaining further information to intercept some of its noise, say $\xi$, and replacing $\hat{\theta}^{\text{cl}}$ by $\hat{\theta}^{\text{cl}}-\xi$ which has MSE $\sigma_{\varepsilon}^{2}$ and remains unbiased for $\theta$. This idea lies at the heart of approaches such as control variates in simulation (Lavenberg & Welch, 1981; Hickernell et al., 2005), variance reduction in randomized controlled experiments via covariate adjustment (Lin, 2013) and by utilizing pre-experiment data (Deng et al., 2013, CUPED), as well as model-assisted estimation in survey sampling (Cochran, 1977; Breidt & Opsomer, 2017). It is also the idea powering PPI and related methods: the unlabeled dataset and the predictive model are used to intercept some of the noise in the classical statistical estimator $\hat{\theta}^{\text{cl}}\triangleq\bar{Y}$; compare to Eq. (2) with $\xi\,\triangleq\,\bar{Z}^{f}-\tilde{Z}^{f}$. We refer to Ji et al. (2025) and Gronsbell et al. (2025) for informative discussions of how PPI relates to traditional ideas in semi-parametric inference as in e.g., Robins et al. (1994).
@@
reduces variance by shrinking $\hat{\theta}^{\text{cl}}$ toward 00 (at the cost of introducing some bias). When $\sigma_{\theta}^{2}$ is small, the MSE of $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}]$ can be substantially smaller than that of $\hat{\theta}^{\text{cl}}$.
Now suppose that the variance of the prior, $\sigma_{\theta}^{2}$, is unknown but we observe data from multiple related problems generated from the same model and indexed by $j\in[m]$, say, $\theta_{j}\,{\buildrel\text{iid}\over{\sim}\,}\mathcal{N}(0,\sigma_{\theta}^{2})$ and $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$. Then an EB analysis can mimic the MSE of the oracle Bayesian that has full knowledge of the prior. To wit, we can estimate $\sigma_{\theta}^{2}$ as
@@
and then consider a plug-in approximation of the Bayes rule, $\hat{\theta}_{j}^{\text{JS}}=\hat{\mathbb{E}}[\theta_{j}\mid\hat{\theta}^{\text{cl}}_{j}]:=\{\hat{\sigma}_{\theta}^{2}/(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\hat{\sigma}_{\theta}^{2})\}\hat{\theta}^{\text{cl}}$. The resulting estimator is the celebrated James-Stein estimator (James & Stein, 1961; Efron & Morris, 1973), whose risk is very close to the Bayes risk under the hierarchical model for large $m$ (Efron, 2010, equation (1.25)). The James-Stein estimator also always dominates the classical estimator under a frequentist evaluation of compound risk in (9) under the assumption that $\hat{\theta}^{\text{cl}}_{j}\,{\buildrel\text{ind}\over{\sim}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$ and $m\geq 3$:
[…]
@@
Instead of using the same prior for each problem, we may try to sharpen the prior and increase its relevance (Efron, 2011) by using further contextual information $\phi$. In the stylized example, as seen in Table 1, such an approach reduces the variance of the prior from $\sigma_{\theta}^{2}$ to $\sigma_{\theta\mid\phi}^{2}<\sigma_{\theta}^{2}$ with corresponding MSE reduction of the Bayes estimator $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}},\phi]$. With multiple related problems, such a strategy can be instantiated via EB shrinkage toward an informative but biased predictor (Fay III & Herriot, 1979; Green & Strawderman, 1991; Mukhopadhyay & Maiti, 2004; Kou & Yang, 2017; Rosenman et al., 2023). The strategy of this form that is closest to our proposal is the covariate-powered EB approach of Ignatiadis & Wager (2019), recently applied to large language model evaluation by Fogliato et al. (2024). Therein (following the notation of Section 2.2), the analyst has access to classical estimators $\bar{Y}_{j}$, $j\in[m]$, and problem-specific covariates $W_{j}$ and seeks to shrink $\bar{Y}_{j}$ toward ML models that predict $\bar{Y}_{j}$ from $W_{j}$. By contrast, in our setting we have observation-level covariates $X_{ij}$ and the ML model operates on these covariates. In principle one could simultaneously use both types of covariates: problem-specific and observation-specific.
@@
One can shrink the variance reduced estimator $\hat{\theta}^{\text{cl}}-\xi$ toward 00 via $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}-\xi]=\{\sigma_{\theta}^{2}/(\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2})\}(\hat{\theta}^{\text{cl}}-\xi)$. In the context of PPI, variance reduction and prior information (with a more heavy-tailed prior) are used by Cortinovis & Caron (2025) within the Frequentist-Assisted by Bayes (FAB) framework of Yu & Hoff (2018)Cortinovis & Caron (2025) only consider a single problem and do not pursue an empirical Bayes approach.

### v3

@@ 1 Introduction
A major obstacle in answering modern scientific questions is the scarcity of gold-standard data (miao2024valid). While advancements in data collection, such as large-scale astronomical surveys (york2000sloan) and web crawling (penedo2024the), have led to an abundance of covariates (or features), scientific conclusions often rely on outcomes (or labels), which are often expensive and labor-intensive to obtain. The rapid development of machine learning (ML) algorithms has offered a path forward, with ML predictions increasingly used to supplement gold-standard outcomes and increase the statistical efficiency of subsequent analyses (liang2007use; wang2020methods).
Prediction-Powered Inference (PPI) (angelopoulos2023prediction) addresses the scarcity issue by providing a framework for valid statistical analysis using predictions from black-box ML models. By combining ML-predicted and gold-standard outcomes, PPI and its variants (angelopoulos_ppi_2024; zrnic2024cross; zrnic2025note) use the abundance of predictions to reduce variance while relying on the accuracy of labeled [^1] Throughout the paper, we use the terms “labeled” and “gold-standard” interchangeably. data to control bias.
In this work, we adapt PPI to the estimation of multiple outcome means in compound estimation settings. Many applications of PPI naturally involve parallel statistical problems that can be solved simultaneously. For instance, several PPI methods (angelopoulos_ppi_2024; fisch2024stratified) have shown improvements in estimating the fraction of spiral galaxies using predictions on images from the Galaxy Zoo 2 dataset (willett2013galaxy). While these methods focus on estimating a single overall fraction, a richer analysis emerges from partitioning galaxies based on metadata (such as celestial coordinates or pre-defined bins) and estimating the fraction of galaxies within each partition. This compound estimation approach enables more granular scientific inquiries that account for heterogeneity across galaxy clusters and spatial locations (nair2010fraction).
We demonstrate, both theoretically and empirically, the benefits of solving multiple mean estimation problems simultaneously. Our approach builds on the empirical Bayes (EB) principle of sharing information across problems (robbins1956empirical; efron2010largescale) as exemplified by James-Stein shrinkage (james1961estimation; xie2012sure). The connection between modern and classical statistical ideas allows us to perform within problem PPI estimation in the first place, followed by a shrinkage step reusing the ML predictions in an adaptive way, which becomes possible through borrowing information across problems. Our contributions are:
@@ 2.1 Prediction-Powered Inference (PPI)
The PPI framework considers a setting where we have access to a small number of labeled data points $(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a large number of unlabeled covariates $(\tilde{X}_{i})_{i=1}^{N}\in(\mathcal{X})^{N}$, where $\mathcal{X}$ and $\mathcal{Y}$ represent the covariate and outcome space, respectively. The data points are drawn iid from a joint distribution $\mathbb{P}_{XY}$. [^2] To be concrete: $(X_{i},Y_{i})\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathbb{P}_{XY}$ and $(\tilde{X}_{i},\tilde{Y}_{i})\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathbb{P}_{XY}$ independently, but $\tilde{Y}_{i}$ is unobserved. We are also given a black-box predictive model $f:\mathcal{X}\to\mathcal{Y}$ that is independent of the datasets (e.g., pre-trained on similar but unseen data). For mean estimation with $\mathcal{Y}\subset\mathbb{R}$, the goal is to leverage the predicted outcomes $f(X_{i})$ to improve the estimation of $\theta:=\mathbb{E}[Y_{i}]$. Some simple estimators take the form of the following aggregated (summary) statistics
$\bar{Y}=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\>\,$ ${\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}\tilde{Y}:=\frac{1}{N}\sum_{i=1}^{N}\tilde{Y}_{i},}$ (1)
$\bar{Z}^{f}=\frac{1}{n}\sum_{i=1}^{n}f(X_{i}),\>\,$ $\tilde{Z}^{f}=\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i}).$
@@
Both definitions represent $\hat{\theta}^{\text{PPI}}$ in the form of a baseline estimator plus a correction term. In the first representation, the baseline estimator is the unbiased classical estimator $\bar{Y}$, while the correction term has expectation 0 and attempts to reduce the variance of $\bar{Y}$. In the second representation, the baseline estimator is the prediction mean on unlabeled data $\tilde{Z}^{f}$ (which in general may be biased for $\theta$), while the correction term removes the bias of $\tilde{Z}^{f}$ by estimating the bias of the ML model $f$ on the labeled dataset. Writing $\hat{\theta}^{\text{PPI}}=\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i})+\tfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))$, we find that $\mathbb{E}[\hat{\theta}^{\text{PPI}}]=\mathbb{E}[Y_{i}]=\theta$ and
@@
To overcome the above limitation, angelopoulos_ppi_2024 introduce a power-tuning parameter $\lambda$ and define
@@
Consider the ideal scenario for PPI with $N=\infty$ (that is, the unlabeled dataset is much larger than the labeled dataset) so that $\tilde{Z}^{f}\equiv\mathbb{E}[f(\tilde{X}_{i})]$. Even then, the MSE of $\hat{\theta}^{\text{PPI}}$ in (3) is always lower bounded [^5] The same lower bound also applies to power-tuned PPI $\hat{\theta}^{\text{PT}}$. by $\tfrac{1}{n}\mathbb{E}[\mathrm{Var}[Y_{i}\mid X_{i}]]$ and the lower bound is attained by the perfect ML predictor $f(\cdot)\equiv\mathbb{E}\!\left[Y_{i}\mid X_{i}=\cdot\right]$. In words, if $Y_{i}$ is not perfectly predictable from $X_{i}$, then PPI applied to a labeled dataset of fixed size $n$ must have non-negligible MSE. By contrast, for $N=\infty$, the prediction mean of unlabeled data $\tilde{Z}^{f}$ has zero variance and MSE equal to the squared bias $(\mathbb{E}[f(X_{i})]-\theta_{i})^{2}$. Thus if the predictor satisfies a calibration-type property that $\mathbb{E}[f(X_{i})]\approx\mathbb{E}[Y_{i}]$ (which is implied by, but much weaker than the requirement $f(X_{i})\approx Y_{i}$), then the MSE of $\tilde{Z}^{f}$ could be nearly 0. By contrast, PPI (and PPI++) can only partially capitalize on such a predictor $f(\cdot)$.
@@ 2.2 The Compound Mean Estimation Setting
In this section, we introduce the problem setting that PAS is designed to address—estimating the mean of $m>1$ parallel problems with a single black-box predictive model ff. [^6] Our proposal also accommodates using separate predictors $\{f_{j}\}_{j=1}^{m}$ for each problem. To streamline exposition, we focus on the practical scenario where a single (large) model (e.g., a large language or vision model) can handle multiple tasks simultaneously (radford2019language; he2022masked). For the $j$-th problem, where $j\in[m]:=\{1,\ldots,m\}$, we observe a labeled dataset $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ samples and an unlabeled dataset $(\tilde{X}_{ij})_{i=1}^{N_{j}}$ with $N_{j}\in\mathbb{N}$ samples. We start with modeling heterogeneity across problems.
@@
$\eta_{j}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathbb{P}_{\eta},\;j\in[m],\,\,\text{and}\,\,\boldsymbol{\eta}:=(\eta_{1},...,\eta_{m})^{\intercal},$ (5)
@@
We do not place any restriction over the unknown prior $\mathbb{P}_{\eta}$. Assumption 2.1 posits exchangeability across problems, which enables information sharing, without restricting heterogeneity (ignatiadis2023empiricala). In our setting, we are specifically interested in the means
@@
[…]
@@
Similar to Eq. (1), we define the aggregated statistics $\bar{Y}_{j},\bar{Z}_{j}^{f},\tilde{Z}^{f}_{j}$ for each $j\in[m]$. To facilitate exposition, following prior work, [^7] For EB, examples include xie2012sure; soloff2024multivariate; for PPI, see recent works like fisch2024stratified. we treat the second moments $\tau_{j}^{2}$, $\sigma_{j}^{2}$, $\gamma_{j}$ as known until the end of Section 5. In Section 6, we extend our methods to allow for unknown $\tau_{j}^{2}$, $\sigma_{j}^{2}$, and $\gamma_{j}$.
@@
$X_{ij}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathcal{N}(\eta_{j},\psi^{2}),\;\;\;Y_{ij}|X_{ij}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{ind}}}\,}\mathcal{N}(2\eta_{j}X_{ij}-\eta_{j}^{2},c).$ (8)
@@
To conclude this section, we define the compound risk (robbins1951asymptotically; jiang2009general) for any estimator $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{m})^{\intercal}$ as the expected squared error loss averaged over problems,
@@ 3 Statistical Guiding Principles & Prior Work
An important statistical idea is to improve $\hat{\theta}^{\text{cl}}$ via obtaining further information to intercept some of its noise, say $\xi$, and replacing $\hat{\theta}^{\text{cl}}$ by $\hat{\theta}^{\text{cl}}-\xi$ which has MSE $\sigma_{\varepsilon}^{2}$ and remains unbiased for $\theta$. This idea lies at the heart of approaches such as control variates in simulation (lavenberg1981perspective; hickernell2005control), variance reduction in randomized controlled experiments via covariate adjustment (lin2013agnostic) and by utilizing pre-experiment data (deng2013improving, CUPED), as well as model-assisted estimation in survey sampling (cochran1977sampling; breidt2017modelassisted). It is also the idea powering PPI and related methods: the unlabeled dataset and the predictive model are used to intercept some of the noise in the classical statistical estimator $\hat{\theta}^{\text{cl}}\triangleq\bar{Y}$; compare to Eq. (2) with $\xi\,\triangleq\,\bar{Z}^{f}-\tilde{Z}^{f}$. We refer to ji2025predictions and gronsbell2025another for informative discussions of how PPI relates to traditional ideas in semi-parametric inference as in e.g., robins1994estimation.
@@
reduces variance by shrinking $\hat{\theta}^{\text{cl}}$ toward 0 (at the cost of introducing some bias). When $\sigma_{\theta}^{2}$ is small, the MSE of $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}]$ can be substantially smaller than that of $\hat{\theta}^{\text{cl}}$.
Now suppose that the variance of the prior, $\sigma_{\theta}^{2}$, is unknown but we observe data from multiple related problems generated from the same model and indexed by $j\in[m]$, say, $\theta_{j}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{iid}}}\,}\mathcal{N}(0,\sigma_{\theta}^{2})$ and $\hat{\theta}^{\text{cl}}_{j}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{ind}}}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$. Then an EB analysis can mimic the MSE of the oracle Bayesian that has full knowledge of the prior. To wit, we can estimate $\sigma_{\theta}^{2}$ as
@@
and then consider a plug-in approximation of the Bayes rule, $\hat{\theta}_{j}^{\text{JS}}=\hat{\mathbb{E}}[\theta_{j}\mid\hat{\theta}^{\text{cl}}_{j}]:=\{\hat{\sigma}_{\theta}^{2}/(\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2}+\hat{\sigma}_{\theta}^{2})\}\hat{\theta}^{\text{cl}}$. The resulting estimator is the celebrated James-Stein estimator (james1961estimation; efron1973stein), whose risk is very close to the Bayes risk under the hierarchical model for large $m$ (efron2010largescale, equation (1.25)). The James-Stein estimator also always dominates the classical estimator under a frequentist evaluation of compound risk in (9) under the assumption that $\hat{\theta}^{\text{cl}}_{j}\,{\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\text{ind}}}\,}\mathcal{N}(\theta_{j},\sigma_{\xi}^{2}+\sigma_{\varepsilon}^{2})$ and $m\geq 3$:
$\mathcal{R}_{m}(\boldsymbol{\hat{\theta}^{\text{JS}}},\boldsymbol{\theta})<\mathcal{R}_{m}(\boldsymbol{\hat{\theta}^{\text{cl}}},\boldsymbol{\theta})\mbox{~~for all~~}\boldsymbol{\theta}\in\mathbb{R}^{m}.$
@@
Instead of using the same prior for each problem, we may try to sharpen the prior and increase its relevance (efron2011tweedie) by using further contextual information $\phi$. In the stylized example, as seen in Table 1, such an approach reduces the variance of the prior from $\sigma_{\theta}^{2}$ to $\sigma_{\theta\mid\phi}^{2}<\sigma_{\theta}^{2}$ with corresponding MSE reduction of the Bayes estimator $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}},\phi]$. With multiple related problems, such a strategy can be instantiated via EB shrinkage toward an informative but biased predictor (fayiii1979estimates; green1991jamesstein; mukhopadhyay2004two; kou2017optimal; rosenman2023combining). The strategy of this form that is closest to our proposal is the covariate-powered EB approach of ignatiadis2019covariatepowered, recently applied to large language model evaluation by fogliato2024precise. Therein (following the notation of Section 2.2), the analyst has access to classical estimators $\bar{Y}_{j}$, $j\in[m]$, and problem-specific covariates $W_{j}$ and seeks to shrink $\bar{Y}_{j}$ toward ML models that predict $\bar{Y}_{j}$ from $W_{j}$. By contrast, in our setting we have observation-level covariates $X_{ij}$ and the ML model operates on these covariates. In principle one could simultaneously use both types of covariates: problem-specific and observation-specific.
@@
One can shrink the variance reduced estimator $\hat{\theta}^{\text{cl}}-\xi$ toward 0 via $\mathbb{E}[\theta\mid\hat{\theta}^{\text{cl}}-\xi]=\{\sigma_{\theta}^{2}/(\sigma_{\varepsilon}^{2}+\sigma_{\theta}^{2})\}(\hat{\theta}^{\text{cl}}-\xi)$. In the context of PPI, variance reduction and prior information (with a more heavy-tailed prior) are used by cortinovis2025fabppi within the Frequentist-Assisted by Bayes (FAB) framework of yu2018adaptivecortinovis2025fabppi only consider a single problem and do not pursue an empirical Bayes approach.